OpenStudy (anonymous):

solve e^2x-3e^x+2=0 for x

OpenStudy (campbell_st):

this is an equation that can be reduced to a quadratic let \[u = e^x\] to start then you have \[u^2 -3u + 2 = 0\] which can be factored \[(u -2)(u-1) = 0\] so u = 2, 1 then reversing the substitution \[e^x = 2..... and..... e^x = 1\] take the ln on both equations to find the value of x

OpenStudy (anonymous):

@campbell_st so would it be ln2 +ln1

OpenStudy (campbell_st):

yep... but ln(1) = 0 so x = 0 or ln(2)

OpenStudy (campbell_st):

and no plus sign between them.

OpenStudy (anonymous):

oh okay :] @campbell_st can you help me with another one? it says solve logx +log (x-1)=log 4x for x.

OpenStudy (campbell_st):

sorry i have to go to school.. post it and there are lots of people here who can help.

OpenStudy (anonymous):

okay well thanks for your help @campbell_st